HH Co. purchased equipment at the beginning of 2020 for $46,000 with an estimated residual value of $4,600 that it has depreciated for the past 2 full years using the double-diminishing-balance method at a rate of 40%. They sold the equipment at the beginning of 2022 for $18,400 cash.

Determine the amount of the gain or loss on the disposal of the equipment?
Answer #1

Depreciation for 2020 = 46,000*40% = 18,400

Depreciation for 2021 = (46,000-18,400)*40% = 11,040

Book value of equipment at the beginning of 2022

= 46,000 - (18,400+11,040)

= 16,560

Gain(Loss) on disposal = Cash received - Book value

= 18,400 - 16,560

Gain of $1840